OpenStudy (anonymous):

Help proving the identity sin^2xcos^2x=(1-cos4x)/8? Medal and fan Thanks(:

OpenStudy (anonymous):

@mathslover could you help me with this one? Unfortunately, this was a huge practice packet and I'm stuck on this question as well.

mathslover (mathslover):

Yes, I will try my best @seehearcreate .

OpenStudy (anonymous):

Awesome, thankyou. I started off with multiplying the squared identities on the left, is that correct?

mathslover (mathslover):

So, we have : \[\sin^2 (x) \cos^2 (x) = \cfrac{(1 - \cos (4x))}{8} \] We will first look at LHS. In the LHS, \(\sin^2 (x) \) can be written as \(1 -\cos^2 (x)\) . Right?

OpenStudy (anonymous):

Right. Would we also write the squared identity of cos^2x?

OpenStudy (anonymous):

By the way, the squared identity of sin^2x is (1-cos^2x)/2, not (1-cos^2)...

OpenStudy (anonymous):

Oh wait, nevermind, I was thinking of something else. Continue.

mathslover (mathslover):

Now, we have : \((1- \cos^2 (x) )(\cos^2 (x)) = \cos^2 (x) - \cos^4 (x) \) Right?

OpenStudy (anonymous):

Yes(:

mathslover (mathslover):

So, now, in RHS you had : \(\cfrac{1 - \cos^4 ( x) }{ 8}\)

OpenStudy (anonymous):

No, it's (1-cos(4x))/8

mathslover (mathslover):

Oh.. Sorry! You need to simplify cos(4x) then.

OpenStudy (anonymous):

So (2cos^2(2x)-1)?

mathslover (mathslover):

Good work.

mathslover (mathslover):

But simplify it further too.

OpenStudy (anonymous):

Okay so 2(2cos^4x-1)-1)?

mathslover (mathslover):

I will get back to you .. have to leave for dinner. May be @ganeshie8 will like to continue from here. Sorry

OpenStudy (anonymous):

Oh no, that's fine, thankyou for your help

OpenStudy (harsha19111999):

We have ... sin² x. cos² x = (1/4) ( 4 sin² x. cos² x ) = (1/4) ( 2 sin x. cos x )² = (1/4) ( sin 2x )² = (1/4) sin² 2x = (1/4) { [ 1 - cos ( 2• 2x ) ] / 2 } = (1/8)( 1 - cos 4x

OpenStudy (anonymous):

Okay, I really don't understand that since you didn't really break it down into distinct steps. If you can add on to what @mathslover was trying to show me, that would be most helpful because that's the work I have down on my sheet.

OpenStudy (anonymous):

@Loser66 Could you help me with this?

OpenStudy (loser66):

@Harsha19111999 gives you the whole thing. Now I just walk you through follow his step, ok? I times both sides by 4 to get \(4 sin^2(x)cos^2(x) = \dfrac{1-cos(4x)}{2}\) OK?

OpenStudy (anonymous):

Alright, but wouldn't it but you would have to have it set up like (1=cos(4x))/2 * (cos^2x) correct?

OpenStudy (loser66):

Be patient and confirm if you get it.

OpenStudy (anonymous):

But you would*

OpenStudy (anonymous):

I do understand that since you multiped the squared sine identity by 4.

OpenStudy (harsha19111999):

Hey didn't you get the answer? Should i explain again?

OpenStudy (loser66):

And we have sin(2x) = 2 sin x cos x so that \((sin (2x))^2 = 4 sin^2(x) cos ^2(x) \) this is our left hand side, right? so that the left hand side can be written as \(sin^2(2x)\) , right?

OpenStudy (anonymous):

No, I really didn't understand it, I needed it broken down into individual steps. Math is my slower subject, unfortunately.

OpenStudy (harsha19111999):

From which step do you want?

OpenStudy (anonymous):

Yes, I understand that.

OpenStudy (anonymous):

The whole thing really. That's why I asked @Loser66 since he breaks down each step until it draws to the conclusion.

OpenStudy (anonymous):

\[\cos 2x=\cos \left( x+x \right)=\cos x~\cos x-\sin x~\sin x=\cos ^2x-\sin ^2 x\] \[=1-\sin ^2x-\sin ^2x=1-2 \sin ^2x \] \[1-\cos 2x=2\sin ^2x\] \[1-\cos 4 x=2 \sin ^2~2x=2\left( 2 \sin x~\cos x \right)^2=8 \sin ^2x~\cos ^2x\] \[\frac{ 1-\cos 4x }{ 8 }=\sin ^2~x \cos ^2~x \]

OpenStudy (anonymous):

Once again, I don't want just the answer, I want the steps explained...I don't understand it otherwise.

OpenStudy (loser66):

Can we continue?? Can you get my stuff?

OpenStudy (anonymous):

I said we could continue..

OpenStudy (loser66):

now, we have the left hand side is \(sin^2(2x) \) right?

OpenStudy (anonymous):

Yes, correct.

OpenStudy (loser66):

I rewrite it: \[\color{red}{sin^2(2x)}= \color {blue}{\dfrac{1-cos (4x)}{2}}\] From the left hand side, I manipulate a little bit. But first off, I want you to know something @joanip_ it's a proving problem, not solving

OpenStudy (anonymous):

That's what @mathslover was doing earlier.

OpenStudy (anonymous):

Right, I understood that ^

OpenStudy (loser66):

\[sin^2 (\color{red}{x}) = \dfrac{1-cos (\color{red}{2x})}{2}\] . This is identity. OK?

OpenStudy (loser66):

and pay attention: the left hand side is x and the right hand side is 2x. OK?

OpenStudy (anonymous):

Yeah, I got that. I know the identities.

OpenStudy (loser66):

so, if the left hand side is 2x, then the right hand side is 4x, right?

OpenStudy (anonymous):

Correct.

OpenStudy (loser66):

\[sin^2(2x)=\dfrac{1+cos(4x)}{2}\] , ok?

OpenStudy (anonymous):

Yep, I understand that completely.

OpenStudy (loser66):

good. hihihihi We are done, right?

OpenStudy (anonymous):

No? the right side was divided by 8?

OpenStudy (anonymous):

When we multiplied the right side by 4, did the 8 reduce down to 2 then?

OpenStudy (loser66):

We manipulate a little bit from the beginning. Re read it

OpenStudy (loser66):

For example: if seehearcreate = Loser 66 and seehearcreate = @dan815 so that, instead of proving seehearcreate = Loser 66, I can prove dan815 = Loser 66 and make conclusion that we are all equal hahahaha

OpenStudy (anonymous):

Yeah, I got you. I forgot about multiplying the right side by 4.

OpenStudy (anonymous):

Thankyou again(:

OpenStudy (loser66):

ok,